OpenStudy (anonymous):

Which can not be written as an equivalent improper fraction? five-thirds three-fourths three and five-sixths twelve-eighths

OpenStudy (anonymous):

Think about it. Do you know what an improper fraction is?

OpenStudy (anonymous):

\[\frac{ 5 }{ 3 } \frac{ 3 }{ 4 } 3\frac{ 5 }{ 6 } or \frac{ 23 }{ 6 }\]\[\frac{ 12 }{ 8 } \] Of all these which one is not an improper fraction?

OpenStudy (anonymous):

thanks @NathanB